OpenStudy (anonymous):

What is 36 ÷ 10^–3? A. 0.0036 B. 0.036 C. 3600 D. 36,000

OpenStudy (anonymous):

10^-3 is the same as 10 raised the third but over 1 so 10^3 = 1 with 3 0's 1000 now put it over 1 1/1000*36 36/1000 this equald B!

jigglypuff314 (jigglypuff314):

@illuminator123 they wanted 36 ÷ 10^–3 you did 36 * 10^-3 :P but it's the same idea :) just 36 / (1/1000) = 36 * 1000 = ?

OpenStudy (anonymous):

i suck thanks for finding that

OpenStudy (anonymous):

thanks everyone Its B